OpenStudy (anonymous):

The cylinder x^2 +y^2 = x divides the unit sphere x^2+ y^2+ z^2=1 into two regions M and N where M is inside the cylinder and N is outside. Compute the ratio of the area of M to the area of N.

OpenStudy (anonymous):

See the attached picture.

OpenStudy (dumbcow):

do you mean ratio of volumes? or surface area? also is there a typo x^2 + y^2 = x is a circle --> (x- .5)^2 +y^2 = .25

OpenStudy (dumbcow):

also, if you can could go here and see if you can help answer this question http://openstudy.com/updates/4fa9f71ce4b059b524f5d875

OpenStudy (anonymous):

Surface areas

OpenStudy (dumbcow):

i get a ratio of: \[\frac{M}{N} = \frac{2\pi -4}{2\pi +4}\] i found area of inside by summing up arc lengths of segments of each circle of unit sphere that was inside cylinder \[\rightarrow 2\int\limits_{0}^{1}r*\cos^{-1} (2r^{2}-1) dS\] \[=2\int\limits_{0}^{1}\frac{r}{\sqrt{1-r^{2}}} \cos^{-1} (2r^{2}-1) dr\] \[= 2\pi - 4\]

OpenStudy (anonymous):

x^2 + y^2 = x is a circle in the plane xy, but a cylinder in 3D where z is a free variable.